OpenStudy (anonymous):

sinx=1/5 and cos y= -12/13, find cos (x+y)

OpenStudy (anonymous):

Here's a trig identity which should help: \[\cos(a+b)=\cos(a)\cos(b)-\sin(a)\sin(b)\]Think you can solve it now?

OpenStudy (anonymous):

I think so but could someone find the final answer to see if I got it right?